# MNRAS  template – title goes here

### v2

@@ 4 Stability Assessment via Dynamical Systems
The techniques of dynamical systems are helpful in examining the overall long-term behavior of a specific cosmological model (Carloni et al., 2005; Amendola et al., 2007; Ivanov & Toporensky, 2012; Carloni et al., 2015; Duchaniya et al., 2023, 2022) and can assist in circumventing the challenge of solving non-linear cosmological equations. These techniques characterize the universal dynamics by studying the nearby long-term behavior of critical points of the system and connecting them to the primary cosmological eras (Paliathanasis, 2017).
@@
In order to study the time-dependent behavior of the dynamical system, it is necessary to establish a dimensionless time parameter. In this study, we choose to use a time parameter expressed as the number of e-folds $N\equiv\text{ln}\,a/a_{0}$, where $a_{0}$ is a constant with the same units as $a$, and is typically set as $a_{0}\equiv 1$. The evolution of each variable is then determined by its derivative with respect to $N$, which is expressed as follows:
@@ 00 $\frac{1}{2}$
  • Critical Point $\mathcal{P}_{2}$: The critical point $\mathcal{P}_{2}$ is depicted in the $XV$ plane in Fig. 8. The critical point $\mathcal{P}_{2}$ corresponds to a non-standard cold dark matter-dominated epoch in which the density of DE is negligible ($\Omega_{\text{DE}}=-20x_{2}$). When $x_{2}=0$, this critical point reflects a standard cold dark matter-dominated era. The phase portrait shows trajectories approaching $\mathcal{P}_{2}$ along specific paths, forming a saddle point structure. This behavior indicates that $\mathcal{P}_{2}$ is a saddle point, with some trajectories being attracted towards it along stable manifolds and repelled away along unstable manifolds. This behavior is consistent with the mixed stability eigenvalues obtained for $\mathcal{P}_{2}$.
  • Critical Point $\mathcal{P}_{3}$: The critical point $\mathcal{P}_{3}$ is illustrated on the $XY$ plane in Fig. 8. It is important to note that the critical point $\mathcal{P}_{3}$, where $q=0$ and $\omega=\frac{-1}{3}$, does not depict accelerating expansion. This critical point occurs at $2m^{2}-m\neq 0$ and is stable when $\frac{1}{4}<m<\frac{1}{2}$. The deceleration parameter solely relies on the variable $Z=\frac{\dot{H}}{H^{2}}$. In this context, when $Z=-1$, it signifies that the contribution from the GB invariant $\mathcal{G}=24H^{4}(Z+1)$ term vanishes, which could explain why it doesn’t demonstrate the transition phase, as depicted in Fig. 8. The phase portrait shows trajectories spiralling inwards towards $\mathcal{P}_{3}$, indicating that it is a stable spiral. This suggests that small perturbations will cause the system to oscillate while eventually converging to $\mathcal{P}_{3}$. The eigenvalues, consisting of negative real parts, confirm the stable nature of this critical point.
  • Critical Point $\mathcal{P}_{4}$: The critical point $\mathcal{P}_{4}$ is depicted on the $XY$ plane in Fig. 8. The critical point $\mathcal{P}_{4}$ represents the de-Sitter solutions with $\Omega_{\text{DE}}=1$, $\Omega_{\text{m}}=0$, $\Omega_{\text{r}}=0$, indicating the current accelerated expansion of the Universe. This de-sitter solution is valid at the critical point $\mathcal{P}_{4}$ within the parameter range $(-1+2m)(-1+4m)\neq 0$ and $m\neq 0$. Consequently, the value of $\omega_{\text{eff}}=q=-1$ highlights the significance of this critical point in describing the current dynamics of the Universe. The trajectories show a clear spiral structure converging towards $\mathcal{P}_{4}$, indicating a stable focus. This implies that the system will exhibit damped oscillations as it approaches $\mathcal{P}_{4}$. The stability analysis of $\mathcal{P}_{4}$ supports this observation, showing that the real parts of the eigenvalues are negative, thus confirming stability.
Understanding the stability and characteristics of critical points is essential for gaining insight into the long-term dynamics and behavior of the specified dynamical system.
@@
In the system described by equations (37), it is possible to perform numerical integration with suitable initial conditions to capture the complete cosmological evolution across different epochs. Current measurements of cosmological parameters (Aghanim et al., 2020) suggest that the Universe is almost flat. For this specific example, we use the initial conditions $X=10^{14}$, $Y=-1.2\times 10^{14}$, $Z=0.005$, and $V=8.2\times 10^{-5}$ and model parameter $m=0.18$. The behavior observed aligns with current cosmic observations regarding the evolution of density parameters. By integrating equation (34) using the summarized initial conditions, we obtain numerical solutions for the density parameters $\Omega_{\text{m}}$, $\Omega_{\text{r}}$, and $\Omega_{\text{DE}}$, as shown in Fig. 9. These results reveal that the Universe evolves through a radiation-dominated phase at early times $(q=1)$. Subsequently, it transitioned into a matter-dominated phase with a deceleration parameter $\frac{1}{2}$. Currently, it is moving into an exponentially accelerating epoch with a deceleration parameter of $-1$. The $f(\mathcal{G})$ model represents a cosmological scenario in which the Universe undergoes successive eras of radiation domination, matter domination, and currently, DE domination. Our results indicate that the point where matter and radiation contribute equally is slightly higher than in the $\Lambda$CDM model. The behavior of the model is consistent with current cosmic observations on the evolution of density parameters. The current densities are approximately $\Omega_{\text{m}}\approx 0.3$, $\Omega_{\text{DE}}\approx 0.7$, and $\Omega_{\text{r}}\approx 10^{-4}$. Similar behavior in the evolution of density parameters have been noted in the literature (Granda (2020)).
@@ 5 Conclusion
In this study, we explored the cosmological properties of a specific modified GB gravity model. Initially, we discussed the main features of a gravitational action, which includes a general combination of the Ricci scalar and the GB invariant. Then, assuming a flat FLRW cosmological background, we derived the point-like Lagrangian of the theory and the corresponding equations of motion. The specific function we focused on, $f(\mathcal{G})=\alpha\sqrt{\beta}(\frac{\mathcal{G}^{2}}{\beta^{2}})^{m}$, approaches GR as the real constant $\alpha$ gets closer to zero. However, our study does not explicitly converge to the cosmological constant case, making it particularly interesting as a potential alternative to the standard $\Lambda$CDM model. This model shows the ability to replicate DE behavior while avoiding the conceptual issues associated with $\Lambda$. Importantly, we showed that the right-hand sides of the modified Friedmann equations can be understood as effective energy density and pressure resulting from curvature.
We investigated the cosmological properties of the $f(\mathcal{G})$ model in the presence of matter fields. We assumed non-relativistic pressureless matter and neglected the late-time contribution of the radiation fluid. By numerically solving the first Friedmann equation, we determined the redshift behavior of the Hubble parameter. We used the $\Lambda$CDM model to establish appropriate initial conditions for $H(z)$ and its derivatives. Subsequently, we utilized the most recent low-redshift observations to compare our theory directly with the model-independent predictions of the cosmic expansion. Specifically, we used a Bayesian analysis with the MCMC method, analysing using the Pantheon+ and CC datasets separately. By assuming uniform prior distributions, we obtained constraints on the free parameters of the model at the 1-$\sigma$ and 2-$\sigma$ confidence levels. This enabled us to reconstruct the cosmological evolution of the Hubble expansion rate and the total effective EoS parameter. Our analysis indicates that the $f(\mathcal{G})$ model effectively accounts for the current acceleration of the Universe without the need for $\Lambda$. However, upon closer analysis and comparison with the predictions of the standard cosmological scenario, it becomes evident that the $f(\mathcal{G})$ model exhibits significant deviations from $\Lambda$CDM as the redshift increases, demonstrating its inability to describe a standard matter-dominated era. Our analysis identifies a notable discrepancy between the $H_{0}$ values obtained from the CC sample and the Pantheon datasets, as seen in Fig. 3, highlighting the ongoing $H_{0}$ tension in cosmology. This result emphasizes the necessity for further research into possible systematic errors or new physics to resolve this issue. Addressing these discrepancies is essential for enhancing our understanding of the expansion rate of the Universe.
@@
In the second phase of our study, we conducted a dynamical system analysis, focusing on the type of $f(\mathcal{G})$ function under consideration. This analysis has enabled us to assess the global behavior and stability of the cosmological model. It provided insights into the critical points associated with the model and their characteristics, which could be relevant to observable cosmology and the evolution of the Universe. Our findings revealed stable critical points describing the late-time cosmic accelerated phase. This indicates non-standard matter and radiation-dominated eras of the Universe. Interestingly, our results align with the standard quintessence model on $z>0$. We made notable preliminary discoveries regarding the finite phase space of a power-law class of the GB gravity model. In addition, the equations describing the dynamical system for the power law $f(\mathcal{G})$ gravity model are provided in Equations (37). Furthermore, Table 3 includes critical points, existing conditions, effective EoS, and deceleration parameters for the autonomous system. In contrast, Table 4 presents density parameter values for the acceleration phase. In total, we identified four critical points, three being stable ($\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$) and one unstable ($\mathcal{P}_{1}$). Finally, Fig. 9 shows the proficiency of the model in depicting the evolution of dark matter, radiation, and DE densities, effectively capturing the transitions through various cosmic epochs and reinforcing the model’s robustness in explaining late-time cosmic phenomena.
@@ Acknowledgments
The authors declare that there is no conflicts of interest.

### v3

@@ 4 Stability Assessment via Dynamical Systems
The techniques of dynamical systems are helpful in examining the overall long-term behaviour of a specific cosmological model (Carloni et al., 2005; Amendola et al., 2007; Ivanov & Toporensky, 2012; Carloni et al., 2015; Duchaniya et al., 2023, 2022) and can assist in circumventing the challenge of solving non-linear cosmological equations. These techniques characterize the universal dynamics by studying the nearby long-term behaviour of critical points of the system and connecting them to the primary cosmological eras (Paliathanasis, 2017).
@@
In order to study the time-dependent behaviour of the dynamical system, it is necessary to establish a dimensionless time parameter. In this study, we choose to use a time parameter expressed as the number of e-folds $N\equiv\text{ln}\,a/a_{0}$, where $a_{0}$ is a constant with the same units as $a$, and is typically set as $a_{0}\equiv 1$. The evolution of each variable is then determined by its derivative with respect to $N$, which is expressed as follows:
@@ 00 $\frac{1}{2}$
  • Critical point $\mathcal{P}_{2}$: The critical point $\mathcal{P}_{2}$ is depicted in the $XV$ plane in Fig. 8. The critical point $\mathcal{P}_{2}$ corresponds to a non-standard CDM-dominated epoch in which the density of DE is negligible ($\Omega_{\text{DE}}=-20x_{2}$). When $x_{2}=0$, this critical point reflects a standard cold dark matter-dominated era. The phase portrait shows trajectories approaching $\mathcal{P}_{2}$ along specific paths, forming a saddle point structure. This behaviour indicates that $\mathcal{P}_{2}$ is a saddle point, with some trajectories being attracted towards it along stable manifolds and repelled away along unstable manifolds. This behaviour is consistent with the mixed stability eigenvalues obtained for $\mathcal{P}_{2}$.
  • Critical point $\mathcal{P}_{3}$: The critical point $\mathcal{P}_{3}$ is illustrated on the $XY$ plane in Fig. 8. It is important to note that the critical point $\mathcal{P}_{3}$, where $q=0$ and $\omega=\frac{-1}{3}$, does not depict accelerating expansion. This critical point occurs at $2m^{2}-m\neq 0$ and is stable when $\frac{1}{4}<m<\frac{1}{2}$. The deceleration parameter solely relies on the variable $Z=\frac{\dot{H}}{H^{2}}$. In this context, when $Z=-1$, it signifies that the contribution from the GB invariant $\mathcal{G}=24H^{4}(Z+1)$ term vanishes, which could explain why it does not demonstrate the transition phase, as depicted in Fig. 8. The phase portrait shows trajectories spiralling inwards towards $\mathcal{P}_{3}$, indicating that it is a stable spiral. This suggests that small perturbations will cause the system to oscillate while eventually converging to $\mathcal{P}_{3}$. The eigenvalues, consisting of negative real parts, confirm the stable nature of this critical point.
  • Critical point $\mathcal{P}_{4}$: The critical point $\mathcal{P}_{4}$ is depicted on the $XY$ plane in Fig. 8. The critical point $\mathcal{P}_{4}$ represents the de Sitter solutions with $\Omega_{\text{DE}}=1$, $\Omega_{\text{m}}=0$, $\Omega_{\text{r}}=0$, indicating the current accelerated expansion of the Universe. This de Sitter solution is valid at the critical point $\mathcal{P}_{4}$ within the parameter range $(-1+2m)(-1+4m)\neq 0$ and $m\neq 0$. Consequently, the value of $\omega_{\text{eff}}=q=-1$ highlights the significance of this critical point in describing the current dynamics of the Universe. The trajectories show a clear spiral structure converging towards $\mathcal{P}_{4}$, indicating a stable focus. This implies that the system will exhibit damped oscillations as it approaches $\mathcal{P}_{4}$. The stability analysis of $\mathcal{P}_{4}$ supports this observation, showing that the real parts of the eigenvalues are negative, thus confirming stability.
Understanding the stability and characteristics of critical points is essential for gaining insight into the long-term dynamics and behaviour of the specified dynamical system.
@@
In the system described by equations (37), it is possible to perform numerical integration with suitable initial conditions to capture the complete cosmological evolution across different epochs. Current measurements of cosmological parameters (Aghanim et al., 2020) suggest that the Universe is almost flat. For this specific example, we use the initial conditions $X=10^{14}$, $Y=-1.2\times 10^{14}$, $Z=0.005$, and $V=8.2\times 10^{-5}$ and model parameter $m=0.18$. The behaviour observed aligns with current cosmic observations regarding the evolution of density parameters. By integrating equation (34) using the summarized initial conditions, we obtain numerical solutions for the density parameters $\Omega_{\text{m}}$, $\Omega_{\text{r}}$, and $\Omega_{\text{DE}}$, as shown in Fig. 9. These results reveal that the Universe evolves through a radiation-dominated phase at early times $(q=1)$. Subsequently, it transitioned into a matter-dominated phase with a deceleration parameter $\frac{1}{2}$. Currently, it is moving into an exponentially accelerating epoch with a deceleration parameter of $-1$. The $f(\mathcal{G})$ model represents a cosmological scenario in which the Universe undergoes successive eras of radiation domination, matter domination, and currently, DE domination. Our results indicate that the point where matter and radiation contribute equally is slightly higher than in the $\Lambda$CDM model. The behaviour of the model is consistent with current cosmic observations on the evolution of density parameters. The current densities are approximately $\Omega_{\text{m}}\approx 0.3$, $\Omega_{\text{DE}}\approx 0.7$, and $\Omega_{\text{r}}\approx 10^{-4}$. Similar behaviour in the evolution of density parameters have been noted in the literature (Granda (2020)).
@@ 5 Conclusion
In this study, we explored the cosmological properties of a specific modified GB gravity model. Initially, we discussed the main features of a gravitational action, which includes a general combination of the Ricci scalar and the GB invariant. Then, assuming a flat FLRW cosmological background, we derived the point-like Lagrangian of the theory and the corresponding equations of motion. The specific function we focused on, $f(\mathcal{G})=\alpha\sqrt{\beta}(\frac{\mathcal{G}^{2}}{\beta^{2}})^{m}$, approaches GR as the real constant $\alpha$ gets closer to zero. However, our study does not explicitly converge to the cosmological constant case, making it particularly interesting as a potential alternative to the standard $\Lambda$CDM model. This model shows the ability to replicate DE behaviour while avoiding the conceptual issues associated with $\Lambda$. Importantly, we showed that the right-hand sides of the modified Friedmann equations can be understood as effective energy density and pressure resulting from curvature.
[…]
@@
In the second phase of our study, we conducted a dynamical system analysis, focusing on the type of $f(\mathcal{G})$ function under consideration. This analysis has enabled us to assess the global behaviour and stability of the cosmological model. It provided insights into the critical points associated with the model and their characteristics, which could be relevant to observable cosmology and the evolution of the Universe. Table 5 presents the eigenvalues of the critical points along with their corresponding stability conditions. Our findings revealed stable critical points describing the late-time cosmic accelerated phase. This indicates non-standard matter and radiation-dominated eras of the Universe. Interestingly, our results align with the standard quintessence model on $z>0$. We made notable preliminary discoveries regarding the finite phase space of a power-law class of the GB gravity model. In addition, the equations describing the dynamical system for the power law $f(\mathcal{G})$ gravity model are provided in Equations (37). Furthermore, Table 3 includes critical points, existing conditions, effective EoS, and deceleration parameters for the autonomous system. In contrast, Table 4 presents density parameter values for the acceleration phase. In total, we identified four critical points, three being stable ($\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$) and one unstable ($\mathcal{P}_{1}$). Finally, Fig. 9 shows the proficiency of the model in depicting the evolution of dark matter, radiation, and DE densities, effectively capturing the transitions through various cosmic epochs and reinforcing the model’s robustness in explaining late-time cosmic phenomena.